OpenStudy (anonymous):

7x-6x+3=14-17 -6+12=3-11+y Solve each.

OpenStudy (anonymous):

x=-6 for the first

OpenStudy (anonymous):

and y=14 just use this http://www.mathpapa.com/algebra-calculator.html

OpenStudy (anonymous):

I LOVE YOU

OpenStudy (cj49):

7x-6x=x x+3=-3 x=-3-3 x=-6